OpenStudy (anonymous):

Simplify 2 square root of 108 6 square root of 3 8 square root of 3 12 square root of 3 72 square root of 3

OpenStudy (anonymous):

Simplify\[2\sqrt{108}\]

OpenStudy (the_fizicx99):

\(\ \sf \Large 2\sqrt{108} \implies 2\sqrt{9 * 12} \implies 6\sqrt{12} \implies 6\sqrt{4*3} = 12\sqrt{3} \)

OpenStudy (abmon98):

o find the square root of this expression... First, look at the "number" part of your expression, the 108 . 108 is 36 times 3 ... ...and 36 can be written as 6 to the power 2. So, the 36 can just be pulled outside of the radical sign, as a 6 , 6\[2*6\sqrt{3}\]